OpenStudy (anonymous):

What is the area of a circle with circumference of 37.68? (Use 3.14 for p.)

OpenStudy (anonymous):

all you have to do is multiply 37.68 and 3.14 to get the area with the circumference

OpenStudy (gorv):

circumference=2*pi*r=37.68 r=37.68/2*pi=6 area =pi*r^2=3.14*6*6